solve for x : 14+[x]=14{x} ; where [.]=greatest integer function and {.}=fractional part

solve for x :
        14+[x]=14{x} ; where [.]=greatest integer function and {.}=fractional part

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
As we know x = I + f
 
hence14 + I + f = 14f
 
14 + I = 13f
 
if I = -14 then 
 
0 = 13 f
f = 0
 
hence x = -14